# Learning from Disengagements: An Analysis of Safety Driver Interventions during Remote Driving

### v1

@@ I Introduction
The remote operation of vehicles is becoming increasingly important, especially in the context of modern mobility concepts and the further development of Automated Driving Systems (ADS). Studies such as the remote driving report published by McKinsey [1] and progress of regulations for remote driving in Germany [2] underline the relevance of the remote driving technology for the future of mobility. However, in order to safely introduce this technology to the market, a well-founded examination of central research questions is required, as emphasized in the Federal Highway Research Institute (BASt) research report for teleoperation needs [3].
@@
However, before specific requirements for the education and training of RDs can be formulated, it is crucial to identify the underlying problems. Previous research has often relied on simulations [7] or controlled test environments [8], often with untrained participants. These approaches provide valuable insights into human behavior under different conditions, but cannot fully reflect the complexity of real-world driving situations [9]. The gap between industrial applications and scientific validation is also highlighted by the increasing number of implementations in an industrial context, for example by companies such as Vay or Fernride. While these companies demonstrate the practical potential of remote driving, scientific research into human performance in real ODDs remains insufficient.
This work addresses the gap in the scientific investigation of Remote Driving Systems (RDS) on public in real-world Operational Design Domains (ODD) and analyses disengagements of the RDS by in-vehicle Safety Driver (SD) interventions. These are situations in which the vehicle control has to be taken over by the in-vehicle SD. The aim is to use this data to draw conclusions about the requirements for training, system safety design and the specification of ODDs in order to create the basis for a safe and effective market introduction of remote driving.
@@ II-A Remote Driving
Remote driving in the automotive industry is distinguished from related concepts such as remote assistance and remote monitoring, forming a nuanced taxonomy of Remote Human Input Systems that varies in complexity [10, 11]. Remote driving involves the direct control [12] of a vehicle by a RD, from a remote location, enabling the navigation of a vehicle through complex environments without requiring physical presence in the vehicle. Building on foundational work by Bogdoll et al. [10] and Amador et al. [11], this field is classified by levels of complexity, which aligns with the SAE taxonomy for ADS [13]. The core of remote driving lies in a stable wireless connection between the Remote Control Station (RCS) and the vehicle, enabling real-time monitoring and control. This connection allows the RD to access the vehicle within ODD remotely, using data transmitted from cameras and further sensors, albeit with latency that impacts real-time the RD’s situational awareness [14, 7].
@@ III Remote Driving System Model
In this work, the RDS developed by Vay Technology was utilized. The RDS consists of three main components: the vehicle, which is equipped with the Vay hardware and software, the RCS, which enables remote control, and the RD as the human-in-the-loop. For the purposes of this work, a SD is used as a additional safety fallback level for RD training purposes. The vehicle was retrofitted to integrate the remote driving technology, including additional cameras and safety controller that monitor and regulate critical safety parameters in real-time. Furthermore, the vehicle’s connectivity was enhanced by specific antennas and modems, integrated with the proprietary Vay connectivity software stack, ensuring a stable and fast communication link between the vehicle and the RCS.
@@
The RCS, shown in Fig. 1, serves as the Human-Machine-Interface (HMI) for the RD, who operates the vehicle remotely. It features three screens that display visual feedback transmitted through multiple camera sensors from the vehicle’s cameras and further sensors, as visualized in Fig. 2. Additionally, microphones and speakers ensure auditory feedback and communication to the inside of the vehicle, while road traffic sound is delivered via external microphones to the RD’s headphone. The vehicle is controlled via a automotive-grade physical steering wheel, along with automotive-grade controls such as column switches, throttle, and brake pedals. Special controllers in the RCS process incoming data and enable interaction with the Vay system installed in the vehicle.
@@ III-A Remote Driving System Human-Machine-Interface
Furthermore, the HMI displays the lateral and longitudinal acceleration and deceleration g-force values based on the current speed. This information supports the RD in stability control in dynamic driving situations such as cornering or braking maneuvers.
@@ III-B Remote Driving System Operational Design Domain
The choice of the ODD ensures that the RDS guarantees the required connectivity and controllability within a defined environment [20]. The ODD of the RDS is specified with regard to the environment, traffic structure, speed limits, weather conditions and other relevant parameters. The RDS used for this study operates exclusively in an urban environment, specifically in Las Vegas, Nevada. The urban environment offers a variety of traffic and infrastructure conditions that the RDS has to cope with. However, the ODD excludes specific parameters such as:
@@ IV Data Preparation and Filtering
  • System Under Test (SUT) exclusion: Data recorded under specific test conditions of the Vay RDS was removed to avoid biases.
@@
Figure 3: Extract of the geographical part of the Remote Driving System (RDS) Operational Design Domain (ODD) and distribution of the remotely driven distance of more than 14,000 km in Las Vegas, Nevada.
@@ V Evaluation
The analysis in this work quantifies the relationship between RD’s driving experience and the number of disengagements and its causes using data collected from a group of 25 RDs. These operated in an ODD, which was designed according to the qualification process of Hans et al. [16]. The RDs, aged between 22 and 44 years ($M=30.44$, $SD=4.81$), where 21 identified as men and 4 identified as women.
All participants completed the standardized remote driving training program from Section III-C in advance. This ensures a uniform level of competence so that the influence of cumulative remote driving experience on driving performance could be investigated without differences in basic skills acting as a confounding factor.
The study covers a total remote-controlled driving distance of 14,291.65 km, spread over several driving sessions within the ODD defined in Section III-B and shown in Fig. 3. During these sessions, performance metrics were continuously collected to analyze the relationship between experience, number of disengagements and disengagement reasons in detail.
RD Level Driving Experience [km] Remotely driven distance [km] Remotely driven duration [mins]
RD–L1 $<$200 4132.56 13,202.53
RD–L2 200–500 5494.06 14,141.65
RD–L3 500–800 4665.04 11,510.48
All Levels 0–800 14,291.65 38854.67
TABLE I: Overview of remotely driven distance and remotely driven minutes by remote driving experience level.
To systematically assess the impact of experience, RDs were categorized into three experience levels based on their cumulative driving distance. Tab. I provides an overview of the remote driving distance and driving duration for each experience level as well as the totals across all levels.
V-A Number of Disengagements over Driving Experience
@@
The curve shows a clear decreasing trend in the number of interventions per 100 km with increasing cumulative remote driving experience. The data suggests that there is a pronounced learning curve with the RDs. Especially in the first 400 km of driving there is a strong reduction in SD interventions. This could be due to a rapid learning of skills or an improved understanding of the RDS. From about 400–500 km, there is a plateau in performance improvement. Additional experience reduces interventions only slightly. This indicates that basic skills are developed at an early stage.
The confidence interval of 90% is relatively wide at lower experience levels, indicating a greater spread in the data. This indicates the need for individualized training strategies to account for different starting levels. With increasing experience, the confidence interval becomes smaller, suggesting a more stable and consistent performance over the group of RDs.
@@
V-B Disengagement Classification
The various categories of disengagements in relation to human RD performance are described below. The classification into these categories creates a basis for the quantitative and qualitative assessment of RD-related disengagements. These are used for the detailed analysis of RD behavior and the identification of specific scenarios that can lead to potential safety-critical situations.
@@ V-B Disengagement Classification
V-C Distribution of Disengagement Reasons
The analysis of the reasons for in total 183 disengagements as a function of the RD’s driving experience provides insights into the causes and frequency of safety-relevant situations during remote driving. A detailed analysis of disengagement reasons is done in Chapter VI and is shown in Fig. 5.
@@ VI Analysis of Disengagement Reasons
In this Section, the reasons listed in Fig. 5 of in total 183 SD interventions are analysed in detail for the causes in order to identify possible reason accumulations or patterns.
@@ VI-A Reason: Braked too late for signs
For RD–L1, 92.59% of theses interventions consisted of situations where the RD recognized the traffic sign (51.85% for traffic lights and 48.15% for stop signs) and verbally called them out, but the braking response was inadequate. This indicates limited adaptation to the long-distance driving environment, possibly due to the partly replaced haptic feedback by visual components, which requires a higher cognitive load and longer adaptation time. This adaptation is more pronounced in experienced RD. In the remaining 7.31% SD interventions of RD–L1, no reaction was observed as the RD was focused on other road users instead of the road sign. Such behavior was no longer observed for RD–L2 and RD–L3, underlining the positive effect of experience on task prioritization and situational awareness.
In terms of responsiveness, RD–L1 showed an insufficient response in 92.6% of scenarios, while 7.4% were classified as a complete absence of response. For RD with more experience than 200 km, there is no longer a lack of response. These differences underline the greater decision-making and perception ability of experienced drivers.
@@ VI-B Reason: Traffic Light went red
SD interventions in which a traffic light changed to red while the remotely driven vehicle was approaching accounted for 7.65% of the total number of disengagements analysed in this study. The number of these incidents remains relatively low across the different experience levels. In this scenarios the speed played a central role, as the majority of 64.3% of these incidents occurred at higher speed within the speed limit, which highlights the challenge of such situations. Specifically, 14.3% incidents occurred at speeds of $<$12 mph, 21.4% incidents occurred at speeds of 12–$<$19 mph and 64.3% incidents occurred at speeds of $\geq 19$ mph. The error rate at higher speeds shows that these situations are particularly challenging, as the short reaction time available and the high speed make decision-making much more difficult. In 71.4% of cases, the reasons for these disengagements are based on the fact that the RD made the wrong decision in the respective situation, either because the RD wanted to cross the traffic light or because the RD applied the brakes too hard.
@@ VI-C Reason: Leaving the lane to the left or right
The analysis of performance in terms of driving experience in remote driving scenarios shows clear differences. Of the 41 scenarios analyzed, 82.9% were performed by RD–L1. While this experience level showed no inadequate reactions in 35.3% of cases, compared to this, there is only one case of RD–L2. This discrepancy indicates a difference in competence, experience and training between these groups.
@@ VII Limitations
First, the results are specific to the RDS and its ODD defined for this study based on a specific use case. While they provide valuable insights, they may not fully translate to other vehicle types such as trucks or other ODDs such as desert environments without further validation. Nevertheless, these results can serve as a starting point for similar applications in other ODDs or vehicle types and provide a basis for further research and adaptation.
@@
Finally, the necessity of the SD interference cannot be clearly answered either. Although there is an mistake by the RD, whether this actually leads to safety-critical scenarios is not answered in this work, as the SD intervention is a precautionary measure.
@@ VIII Discussion
The first scenario, shown in Fig. 6(a), at a speed of $\geq 6$ mph where the RD reacts but does not manage to stop in time in front of a sign. This maneuver accounts for 18.03% of all disengagements, of which 13.66% can be attributed to RDs with a driving experience of $<$200 km. This means that this event occurs every 165.30 km for the RD-L1 and only every 1269.89 km for the RD–L2 and RD–L3.
@@
Leaving the lane during turning left or right at a speed of $<$12 mph is responsible for a total of 12.57% of the disengagements and represents scenario 3, which is shown in Fig. 6(c). Here, the RD-L1 are responsible for a full 11.48%, which means that the RD–L1 experience this scenario every 196.79km, while RD with a remote driving experience of 200–800 km only experience this scenario every 5079.55 km. The reason for this is possibly the additional latency, the insufficient haptic feedback or distorted spatial perception.
The proportion of the scenario, visualised in Fig. 6(d), where a traffic light changes to yellow/red at a speed of $\geq 12$ mph is 6.01%. Here the RD reacted, but according to the assessment of the SD for wrong and the SD had to take over based on the respective situation.
@@
Future work should investigate whether the identified problems and specific scenarios described in this work can already be assessed using other driving metrics, so that constructive or instructive measures can be taken at an early stage in order to prevent SD intervention.

### v2

@@ I Introduction
The remote operation of vehicles is becoming increasingly important, especially in the context of modern mobility concepts and the further development of Automated Driving Systems (ADS). Studies such as the remote driving report published by McKinsey [1] and progress of regulations for remote driving in Germany [2] underline the relevance of the remote driving technology for the future of mobility.
However, in order to safely introduce this technology to the market, a well-founded examination of central research questions is required, as emphasized in the Federal Highway Research Institute (BASt) research report for teleoperation needs [3].
@@
However, before specific requirements for the education and training of RDs can be formulated, it is crucial to identify the underlying problems. Previous research has often relied on simulations [7] or controlled test environments [8], often with untrained participants. These approaches provide valuable insights into human behavior under different conditions, but cannot fully reflect the complexity of real-world driving situations [9].
The gap between industrial applications and scientific validation is also highlighted by the increasing number of implementations in an industrial context, for example by companies such as Vay Technology or Fernride. While these companies demonstrate the practical potential of remote driving, scientific research into human performance in real-world Operational Design Domains (ODD) remains insufficient.
This work addresses the gap in the scientific investigation of Remote Driving Systems (RDS) on public in real-world ODD and analyses disengagements of the RDS by in-vehicle Safety Driver (SD) interventions. These are situations in which the vehicle control has to be taken over by the in-vehicle SD to prevent a potential hazardous situations. The aim is to use this data to draw conclusions about the requirements for training, system safety design and the specification of ODDs in order to create the basis for a safe and effective market introduction of remote driving.
@@ II-A Remote Driving
Remote driving in the automotive industry is distinguished from related concepts such as remote assistance and remote monitoring, forming a nuanced taxonomy of Remote Human Input Systems that varies in complexity [10]. Remote driving involves the direct control [11] of a vehicle by a RD, from a remote location, enabling the navigation of a vehicle through complex environments without requiring physical presence in the vehicle. Building on foundational work by Bogdoll et al. [10] and Amador et al. [12], this field is classified by levels of complexity, which aligns with the SAE taxonomy for ADS [13]. The core of remote driving lies in a stable wireless connection between the Remote Control Station (RCS) and the vehicle, enabling real-time monitoring and control. This connection allows the RD to access the vehicle within ODD remotely, using data transmitted from cameras and further sensors, albeit with latency that impacts real-time the RD’s situational awareness [14, 7].
@@ III Remote Driving System Model
In this work, the RDS developed by Vay Technology was utilized. The RDS consists of three main components: the vehicle, which is equipped with the Vay hardware and software, the RCS, which enables remote control, and the RD as the human-in-the-loop. For the purposes of this work, a SD is used as a additional safety fallback level for RD training purposes. The RD acts as the vehicle driver and communicates any driving intentions or changes in the immediate surroundings. This enables the SD to understand the RD’s intended actions. However, if the RD deviates from the announced intention, commits a traffic violation or creates a potentially unsafe situation, the SD takes over control.
The vehicle was retrofitted to integrate the remote driving technology, including additional cameras and safety controller that monitor and regulate critical safety parameters in real-time. Furthermore, the vehicle’s connectivity was enhanced by specific antennas and modems, integrated with the proprietary Vay connectivity software stack, ensuring a stable and fast communication link between the vehicle and the RCS.
@@
The RCS, shown in Fig. 1, serves as the Human-Machine-Interface (HMI) for the RD, who operates the vehicle remotely. It features three screens that display visual feedback transmitted through multiple camera sensors from the vehicle’s cameras and further sensors, as visualized in Fig. 2. Additionally, microphones and speakers ensure auditory feedback and communication to the inside of the vehicle, while road traffic sound is delivered via external microphones to the RD’s headphone. To make it more realistic for RDs to control a vehicle, a concept of torque emulation is used, introduced by Hossain et al. [21]. In addition to that, steering force feedback is provided to the RD. The vehicle is controlled via a automotive-grade physical steering wheel, along with automotive-grade controls such as column switches, throttle, and brake pedals. Special controllers in the RCS process incoming data and enable interaction with the Vay system installed in the vehicle.
@@ III-A Remote Driving System Human-Machine-Interface
Furthermore, the HMI displays the lateral and longitudinal acceleration and deceleration g-force values based on the current speed change. This information supports the RD in stability control in dynamic driving situations such as cornering or braking maneuvers.
@@ III-B Remote Driving System Operational Design Domain
The definition of the ODD ensures that the RDS guarantees the required connectivity and controllability within a defined environment [22]. The ODD of the RDS is specified with regard to the environment, traffic structure, speed limits, weather conditions and other relevant parameters. The RDS used for this study operates exclusively in an urban environment, specifically in Las Vegas, Nevada. The urban environment offers a variety of traffic and infrastructure conditions that the RDS has to cope with. However, the ODD excludes specific parameters such as:
@@ III-C Remote Driver Training
RD level Cumulative driving Experience [km] Remotely driven distance [km] Remotely driven duration [mins] Average speed [mph] Average speed [km/h] RD contributed
RD–L1 $<$200 4,132.56 13,202.53 11.16 17.96 25
RD–L2 200–500 5,494.06 14,141.65 14.03 22.58 18
RD–L3 500–800 4,665.04 11,510.48 14.35 23.09 16
All levels 0–800 14,291.65 38,854.67 13.02 20.96 25
TABLE I: Overview of remotely driven distance, remotely driven duration, and average speed across Remote Driver (RD) experience levels.
@@ IV Data Preparation and Filtering
  • System Under Test exclusion: Data recorded under specific test conditions of the Vay RDS was removed to avoid biases.
@@
Figure 3: Extract of the geographical part of the Remote Driving System (RDS) Operational Design Domain (ODD) and distribution of the remotely driven distance of more than 14,000 km in Las Vegas, Nevada, US.
@@ V Evaluation
The analysis in this work quantifies the relationship between RD’s driving experience and the number of disengagements and its causes using data collected from a group of 25 RDs. These operated in an ODD, which was designed according to the ODD qualification process of Hans et al. [17]. The RDs were aged between 22 and 44 years ($M=30.44$, $SD=4.81$), where 21 identified as men and 4 identified as women. All participants completed the standardized remote driving training program from Section III-C in advance. This ensures a uniform level of competence so that the influence of cumulative remote driving experience on driving performance could be investigated without differences in basic skills acting as a confounding factor.
The study covers a total driving distance of 14,291.65 km, spread over several remote driving sessions within the ODD defined in Section III-B and shown in Fig. 3. During these sessions, performance metrics were continuously collected to analyze the relationship between cumulative remote driving experience, number of disengagements and their reasons in detail. To systematically assess the impact of various influencing factors on remote driving performance, the data analysis considered the three dimensions (1) the driving speed, (2) the level of driving experience, and (3) the driving scenario type. RDs were first categorized into three experience levels based on their cumulative remotely driven distance. Table I provides an overview of the total remotely driven distance and driving duration for each experience level, as well as aggregated values across all levels. In addition to this, each driving event was classified by scenario type (e.g., intersection vs. straight road segments) and by speed range (slow, medium, or high), based on predefined thresholds. This multi-factorial categorization enabled a differentiated analysis of performance-related indicators across varying operational contexts.
V-A Evaluation of average driving speed
The average driving speed per remotely driven session increases as shown in Table I, indicating a relationship between experience and operational efficiency. RD–L1 drivers exhibit a lower average speed of 11.16 mph, compared to 14.03 mph for RD–L2 and 14.35 mph for RD–L3 drivers. This difference reflects growing driving competence and possibly greater trust either in the RDS or in one’s own abilities as experience increases. This is confirmed by a statistically significant, moderately positive correlation between certification level and average speed. Specifically, a Pearson correlation of $r=0.446$ ($p<0.001$) and a Spearman correlation of $\rho=0.452$ ($p<0.001$) suggest that more experienced RDs tend to operate at higher average speeds. At the same time, the reduced speed observed that RD–L1 may reflect a more cautious driving style typical of the early stages of driving experience. The small difference in average speed per session between RD–L2 and RD–L3 suggests that the most substantial gain in driving competence occurs during the transition from RD–L1 to RD–L2, while further experience beyond this point has a diminishing impact on average speed.
V-B Number of Disengagements over Driving Experience
@@
The curve shows a clear decreasing trend in the number of interventions per 100 km with increasing cumulative remote driving experience. The data suggests that there is a pronounced learning curve with the RDs. Especially in the first 400 km of driving there is a strong reduction in SD interventions. This could be due to a rapid learning of skills or an improved understanding of the RDS. From about 400–500 km, there is a plateau in performance improvement. Additional experience reduces interventions only slightly. This indicates that basic skills are developed at an early stage. The confidence interval of 90% is relatively wide at lower experience levels, indicating a greater spread in the data. This indicates the need for individualized training strategies to account for different starting levels. With increasing experience, the confidence interval becomes smaller, suggesting a more stable and consistent performance over the group of RDs.
@@
V-C Disengagement Classification
The various categories of disengagements in relation to human RD performance are described below. The classification is carried out post-hoc on the basis of video material recorded during the remote drives. The situations in which the SD had to intervene in order to avoid an unsafe or irregular driving situation were analyzed. The classification into these categories creates a basis for the quantitative and qualitative assessment of RD-related disengagements. These are used for the detailed analysis of RD behavior and the identification of specific scenarios that can lead to potential safety-critical situations.
@@ V-C Disengagement Classification
V-D Distribution of Disengagement Reasons
The analysis of the reasons for in total 183 disengagements as a function of the RD’s cumulative driving experience provides insights into the causes and frequency of safety-relevant situations during remote driving. A detailed analysis of disengagement reasons is done in Chapter VI and is shown in Fig. 5.
@@ VI Analysis of Disengagement Reasons
In this Section, the reasons listed in Fig. 5 of in total 183 SD interventions are analyzed in detail for the causes in order to identify possible reason accumulations or patterns.
@@ VI-A Reason: Braked too late for signs
For RD–L1, 92.59% of theses interventions consisted of situations where the RD recognized the traffic sign (51.85% for traffic lights and 48.15% for stop signs) and verbally called them out, but the braking response was inadequate. This indicates limited adaptation to the long-distance driving environment, possibly due to the partly replaced haptic feedback by visual components, which requires a higher cognitive load and longer adaptation time. This adaptation is more pronounced in experienced RD. In the remaining 7.31% SD interventions of RD–L1, no reaction was observed as the RD was focused on other road users instead of the road sign. Such behavior was no longer observed for RD–L2 and RD–L3, underlining the positive effect of experience on task prioritization and situational awareness. These differences underline the greater decision-making and perception ability of experienced drivers.
@@ VI-B Reason: Traffic Light went red
SD interventions in which a traffic light changed to red while the remotely driven vehicle was approaching accounted for 7.65% of the total number of disengagements analyzed in this study. The number of these incidents remains relatively low across the different experience levels. In this scenarios the speed played a central role, as 14.3% incidents occurred at speeds of $<$12 mph, 21.4% incidents occurred at speeds of 12–$<$19 mph and 64.3% incidents occurred at speeds of $\geq 19$ mph. The error rate at higher speeds shows that these situations are particularly challenging, as the short reaction time available and the high speed make decision-making much more difficult. In 71.4% of cases, the reasons for these disengagements are based on the fact that the RD made the wrong decision in the respective situation, either because the RD wanted to cross the traffic light or because the RD applied the brakes too hard. In the other 28.6% of cases, the RD responded too late, too slowly or not at all.
@@ VI-C Reason: Leaving the lane to the left or right
The analysis of performance in terms of driving experience in remote driving scenarios shows clear differences. Of the 41 scenarios analyzed, 82.9% were performed by RD–L1. While this experience level showed no inappropriate reactions in 35.3% of cases, there is only one case of RD–L2 in comparison, which indicates a higher competence, experience and training for more experienced RDs.
@@ VII Limitations
First, the results are specific to the RDS and its ODD defined for this study based on a specific use case. While they provide valuable insights, they may not fully translate to other vehicle types such as trucks or other ODDs such as desert environments without further validation. Nevertheless, these results can serve as a starting point for similar applications in other ODDs or vehicle types and provide a basis for further research.
@@
Finally, the necessity of the SD interference as a precautionary measure cannot be clearly answered. Although there is an mistake by the RD, whether this actually leads to safety-critical scenarios is not answered in this work.
@@ VIII Discussion
The analysis focuses on the disengagements by the RD, which are primarily due to human perception and decision-making errors, while in the case of ADS they result primarily from technical limitations, especially in the area of environment recognition and decision logic under uncertainty [18]. While RD can demonstrably reduce its susceptibility to errors through targeted training, ADS requires structural improvements to AI architectures and more robust behavior in complex traffic scenarios. The results make it clear that the causes of disengagements depend heavily on the system type and that differentiated optimization strategies are required accordingly.
@@
The first scenario, shown in Fig. 6(a), at a speed of $\geq 6$ mph where the RD reacts but does not manage to stop in time in front of a sign. This maneuver accounts for 18.03% of all disengagements, of which 13.66% can be attributed to RDs with a driving experience of $<$200 km. This means that this event occurs every 165.30 km for the RD-L1 and only every 1,269.89 km for the RD–L2 and RD–L3.
@@
Leaving the lane during turning left or right at a speed of $<$12 mph is responsible for a total of 12.57% of the disengagements and represents scenario 3, which is shown in Fig. 6(c). Here, the RD-L1 are responsible for a full 11.48%, which means that the RD–L1 experience this scenario every 196.79km, while RD with a remote driving experience of 200–800 km only experience this scenario every 5,079.55 km. The reason for this is possibly the additional latency, the insufficient haptic feedback or distorted spatial perception.
The proportion of the scenario, visualized in Fig. 6(d), where a traffic light changes to yellow/red at a speed of $\geq 12$ mph is 6.01%. Here the RD reacted, but according to the assessment of the SD for wrong and the SD had to take over based on the respective situation.
@@
Future work should investigate whether the identified problems and specific scenarios described in this work can already be assessed with other driving metrics, so that constructive or instructive measures can be taken at an early stage to prevent SD interventions. In addition, possible measures to improve training should be discussed, such as the targeted use of simulator training to systematically prepare RDs for typical error scenarios and improve their situational decision-making behavior. At the same time, it remains unclear whether the RD errors analyzed in the study would actually have led to safety-critical situations if no intervention had taken place. Although the categorization of disengagement is based on observable driving errors, a clear causal relationship between RD behavior and real hazard potential is not conclusively established. Future research should therefore differentiate more strongly between actual and potential hazards.